HETAV PATEL

Grade 11,

If cosx+cosy-cos(x+y)=3/2 then find the value of SinxSiny where x>0, y

If cosx+cosy-cos(x+y)=3/2 then find the value of SinxSiny where x>0, y
 

Grade:11

1 Answers

noogler
8 years ago
 
cosx+cosy-cos(x+y)=3/2 is true if x=y=60o
cos60o+cos60o-cos120o=1/2+1/2-(-1/2)=3/2
SinxSiny=sin60o sin60o=3/4
sinx siny=3/4
 
approve if it is r8 and useful to u..........